OpenStudy (anonymous):

Find all z ∈ ℂ which satisfy z^2 + 2z + 1 = 0.

OpenStudy (anonymous):

I need help with the working. I am sure if this is the correct way, but I reduced the equation to \[a ^{2} + 2a + 1 +(2a + bi + 2)bi = 0\]

OpenStudy (anonymous):

i am not sure

OpenStudy (anonymous):

but from there i'm not sure how to proceed

OpenStudy (ash2326):

instead of putting z=a+bi, you just find first the value of z equation is z^2+z+1=0 so (z+1)^2=0 z=-1 so z is purely real if we write in terms of a+bi then here it's -1+0i the other root is -1-0i so both the roots are same here,and purely real

OpenStudy (anonymous):

Thx! :)

OpenStudy (ash2326):

glad to help